OpenStudy (anonymous):

Please help I stink at graphs. Will fan and medal! Choose the graph below that represents the following system of inequalities: y greater than or equal to -3x + 1 y greater than or equal to 1 over 2x + 3 Graph of two lines that intersect at one point. Both lines are solid. One line passes through points (-2, 2) and (0, 3) and is shaded above the line. The other line passes through points (0, 1) and (1, -2) and is shaded above the line. Graph of two lines that intersect at one point. Both lines are solid. One line passes through points (-2, 2) and (0, 3) and is shaded below the

OpenStudy (igreen):

Okay, we can graph this and see which description best fits it. \(y \ge -3x + 1\) \(y \ge \dfrac{1}{2}x + 3\) Here's the graph: https://www.desmos.com/calculator/xgkhthojvs

OpenStudy (igreen):

Which do you think is correct?

OpenStudy (igreen):

And I don't think you included all of your Options.

OpenStudy (anonymous):

Graph of two lines that intersect at one point. Both lines are solid. One line passes through points (-2, 2) and (0, 3) and is shaded above the line. The other line passes through points (0, 1) and (1, -2) and is shaded above the line. Graph of two lines that intersect at one point. Both lines are solid. One line passes through points (-2, 2) and (0, 3) and is shaded below the line. The other line passes through points (0, 1) and (1, -2) and is shaded above the line. Graph of two lines that intersect at one point. Both lines are solid. One line passes through points (-2, 2) and (0, 3) and is shaded below the line. The other line passes through points (0, 1) and (1, -2) and is shaded below the line. Graph of two lines that intersect at one point. Both lines are solid. One line passes through points (-2, 2) and (0, 3) and is shaded above the line. The other line passes through points (0, 1) and (1, -2) and is shaded below the line.

OpenStudy (anonymous):

I was thinking it was c but I'm so terrible at graphs it could be anything

OpenStudy (igreen):

This one is better: https://www.desmos.com/calculator/aav3uw4yop

OpenStudy (igreen):

Option C says that they are both shaded below the line..but that is not true, both of them are actually shaded above the line.

OpenStudy (anonymous):

I'm going to be honest no matter how any times someone explains them to me I never understand graphs

OpenStudy (igreen):

Lol, the key here is if the lines are shaded above or below the line, because they are different for every option. See which option says that BOTH of the two lines are shaded above the line.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so it's a but would you mind explaining in more depth how to get that answer

OpenStudy (igreen):

Okay, yes, A is correct. It tells you what points they pass through, you can confirm this by plugging them into the inequalities and see if they're true (just like the last question we did), it also tells you about the line being solid or dotted. Solid lines are used when we have a less than or equal to sign(\(\le\)) or a greater than or equal to sign(\(\ge\)). Dotted lines are used when we have a less than sign(\(<\)) or a greater than sign (\(>\)). If it is shaded above the line, the inequality sign will be \(\ge\) or \(>\), but if it is shaded below the line, the inequality sign will be \(\le\) or \(<\). So we don't have to graph it at all to see what the line's will look like.

OpenStudy (igreen):

So basically it's all about the inequality they use.

OpenStudy (anonymous):

Thank you so much

OpenStudy (igreen):

No problem.